# Semiclassical Gravity Beyond General Relativity: Insights from Torsion

### v1

@@ 3.3.2 Conformal transformations
$g^{ab}\rightarrow\Omega^{-2}g^{ab},\quad K^{c}_{\ ab}\rightarrow K^{c}_{\ ab},\quad\Phi\rightarrow\Omega^{-1}\Phi,$
@@
$\delta S_{\text{M}}=\int\epsilon\,\tau^{a}_{\ a}\,\sqrt{-g}\,\mathrm{d}^{4}x.$ (36)
@@
$\tau^{a}_{\ a}=0.$ (37)
For the Klein–Gordon field under consideration, and in the special case where $m=0$, $\xi=1/6$, and the fields are on-shell, this reduces to
$\tau^{a}_{\ a}=\frac{1}{3}\mathring{\nabla}_{a}\big(K^{a\ b}_{\ b}\,\Phi^{2}\big),$ (38)
@@ 4 Hadamard bi-parametrix
To obtain well-defined operators associated with $\tau_{ab}$ and $\sigma_{c}^{\ ab}$, which are quadratic in $\hat{\Phi}$, we employ point-splitting renormalization [38]. In this scheme, the central object is the two-point function
@@
We now insert Eq. (4) into Eq. (4) for $x\neq x^{\prime}$. From Eqs. (107) and (117) in the Appendix, we obtain
@@
where $\Delta$ is the Van Vleck–Morette determinant (see the Appendix). All arguments are omitted for brevity. By collecting the terms with the same $\sigma$-dependence, we can extract the following relations:
@@
From Eqs. (55) and (56), we can verify that $q_{a_{1}\cdots a_{p}}=q_{(a_{1}\cdots a_{p})}$, where $q$ stands for either $v_{n}$ or $w_{n}$. This symmetry, together with the exchange symmetry $x\leftrightarrow x^{\prime}$ of $V_{0}$ and $W$, allows us to express the coefficients with an odd number of indices in terms of those with an even number of indices. This is achieved by using Eqs. (108)-(115) given in Appendix 7, explicitly implementing the exchange symmetry in Eqs. (55) and (56), and then taking the coincidence limit. The resulting identities are
@@
$w^{a}_{\ a}$ $=(m^{2}+\xi\mathring{R})w-6v_{1}+2\xi(\mathring{\nabla}_{a}K^{a\ b}_{\ b}+K^{c\ a}_{\ [a}K^{b}_{\ b]c})w,$ (60)
@@
[…]
${v_{0}}^{a}_{\ a}$ $=(m^{2}+\xi\mathring{R})v_{0}-4v_{1}+2\xi(\mathring{\nabla}_{a}K^{a\ b}_{\ b}+K^{c\ a}_{\ [a}K^{b}_{\ b]c})v_{0},$ (62)
$\mathring{\nabla}_{b}{v_{0}}^{b}_{\ a}$ $=\frac{1}{4}\mathring{\nabla}_{a}\square v_{0}+\frac{1}{2}\mathring{R}_{ab}\mathring{\nabla}^{b}v_{0}+\frac{1}{2}\xi\mathring{\nabla}_{a}\mathring{R}v_{0}-\mathring{\nabla}_{a}v_{1}$
@@
$+\frac{1}{90}\mathring{R}_{a}^{\ c}\mathring{R}_{bc}-\frac{1}{180}\mathring{R}^{cd}\mathring{R}_{acbd}-\frac{1}{180}\mathring{R}_{a}^{\ cde}\mathring{R}_{bcde},$ (67)
$v_{1}$ $=\frac{1}{8}m^{4}+\frac{1}{4}\xi m^{2}R-\frac{1}{24}m^{2}\mathring{R}-\frac{1}{24}\xi\square R+\frac{1}{120}\square\mathring{R}+\frac{1}{8}\Big(\xi R-\frac{1}{6}\mathring{R}\Big)^{2}$
$-\frac{1}{720}\mathring{R}_{ab}\mathring{R}^{ab}+\frac{1}{720}\mathring{R}_{abcd}\mathring{R}^{abcd}.$ (68)
@@ 5 Semiclassical Einstein–Cartan theory
$\mathring{G}_{ab}+K^{d}_{\ (ab)}K^{c}_{\ cd}-K^{d}_{\ c(a}K^{c}_{\ b)d}$ $-K^{d\ \ c}_{\ [c}K^{e}_{\ e]d}g_{ab}$
@@
$g^{ab}K^{d}_{\ dc}-\delta^{a}_{\ c}K^{d\ b}_{\ d}+K_{c}^{\ ba}-K^{b\ a}_{\ c}$ $=16\pi\,\omega(\sigma_{c}^{\ ab}),$ (70)
@@ 5.1 Axiomatic framework
    is smooth for any pair of test functions of compact support, $f_{1}$, $f_{2}$, then the differences $\omega_{1}(\tau_{ab})-\omega_{2}(\tau_{ab})$ and $\omega_{1}(\sigma_{c}^{\ ab})-\omega_{2}(\sigma_{c}^{\ ab})$ are also smooth.
@@
    The expectation values $\omega(\tau_{ab}(x))$ and $\omega(\sigma_{c}^{\ ab}(x))$ are local with respect to the state of the Klein–Gordon field in the following sense: Let $(\mathcal{M},g_{ab})$ and $(\mathcal{M}^{\prime},g_{ab}^{\prime})$ be globally hyperbolic spacetimes with Cauchy surfaces $\Sigma$ and $\Sigma^{\prime}$, respectively. Let $\mathcal{O}\subset\mathcal{M}$ and $\mathcal{O}^{\prime}\subset\mathcal{M}^{\prime}$ be globally hyperbolic open neighborhoods of $x$ and $x^{\prime}$, with Cauchy surfaces $\mathcal{O}\cap\Sigma$ and $\mathcal{O}^{\prime}\cap\Sigma^{\prime}$, such that there exists an isometry between $\mathcal{O}$ and $\mathcal{O}^{\prime}$. Under this isometry, we can identify the Klein–Gordon subalgebras ${\mathcal{A}_{\text{KG}}}_{\mathcal{O}}\subset\mathcal{A}_{\text{KG}}$ for $\mathcal{M}$ and ${\mathcal{A}^{\prime}_{\text{KG}}}_{\mathcal{O}^{\prime}}\subset\mathcal{A}^{\prime}_{\text{KG}}$ for $\mathcal{M}^{\prime}$. If the restrictions of the Hadamard states coincide, i.e. $\omega|_{{\mathcal{A}_{\text{KG}}}_{\mathcal{O}}}=\omega^{\prime}|_{{\mathcal{A}^{\prime}_{\text{KG}}}_{\mathcal{O}^{\prime}}}$, then we require that $\omega(\tau_{ab}(x))=\omega^{\prime}(\tau_{ab}(x^{\prime}))$ and $\omega(\sigma_{c}^{\ ab}(x))=\omega^{\prime}(\sigma_{c}^{\ ab}(x^{\prime}))$.
@@
    $(K^{c}_{\ bd}\mathring{\nabla}_{a}+K^{c}_{\ ab}\mathring{\nabla}_{d}+\mathring{\nabla}_{a}K^{c}_{\ bd}+\mathring{\nabla}_{d}K^{c}_{\ ab})\,\omega(\sigma_{c}^{\ ad})$
    $-(K^{d}_{\ ac}\mathring{\nabla}_{d}+\mathring{\nabla}_{d}K^{d}_{\ ac})\,\omega(\sigma_{b}^{\ ac})-\mathring{\nabla}_{b}K^{c}_{\ ad}\,\omega(\sigma_{c}^{\ ad}).$ (72)
@@
    For a vacuum $\Omega$ in flat, torsionless spacetime, $\omega_{\Omega}(\tau_{ab})=0$ and $\omega_{\Omega}(\sigma_{c}^{\ ab})=0$.
The first two axioms determine $\omega(\tau_{ab})$ and $\omega(\sigma_{c}^{\ ab})$ up to local curvature terms, which cannot diverge in the flat-spacetime limit, as required by the fourth axiom. The third axiom is directly inspired by the (classical) relation given in Eq. (32). Moreover, the fourth axiom ensures the consistency of the semiclassical approach with quantum field theory in torsion-free flat spacetime. Equipped with these axioms, we now perform the regularization of the expectation values of the energy–momentum and spin–density operators.
@@ 5.2 Renormalization
$-K^{c}_{\ (ab)}\mathring{\nabla}_{c}-K^{c}_{\ (ab)}g_{c}^{\ c^{\prime}}\mathring{\nabla}_{c^{\prime}}$
$+K^{d}_{\ (ab)}K^{c}_{\ cd}-K^{d}_{\ c(a}K^{c}_{\ b)d}$
$-g_{ab}(K^{d\ \ c}_{\ [c}K^{e}_{\ e]d}-K^{c\ d}_{\ d}\mathring{\nabla}_{c}-K^{c\ d}_{\ d}g_{c}^{\ c^{\prime}}\mathring{\nabla}_{c^{\prime}})],$ (73)
@@
$+g^{ab}K^{d}_{\ dc}-\delta^{a}_{\ c}K^{d\ b}_{\ d}+K_{c}^{\ ba}-K^{b\ a}_{\ c}),$ (74)
where unprimed indices refer to $x$, and primed indices to $x^{\prime}$. In addition, $g_{a}^{\ a^{\prime}}$ is the bitensor that implements parallel transport from $x$ to $x^{\prime}$, as described in the Appendix. Concretely, the expectation values we need to compute are
@@
$\omega(\sigma_{c}^{\ ab})$ $=\lim_{x^{\prime}\rightarrow x}\hat{\sigma}_{c}^{\ ab}(x,x^{\prime})[-iG^{\text{F}}(x,x^{\prime})].$ (76)
@@
$\omega(\sigma_{c}^{\ ab})$ $=\lim_{x^{\prime}\rightarrow x}\hat{\sigma}_{c}^{\ ab}(x,x^{\prime})[-i(G^{\text{F}}(x,x^{\prime})-H_{\ell}(x,x^{\prime}))]$
$=\frac{1}{8\pi^{2}}\lim_{x^{\prime}\rightarrow x}\hat{\sigma}_{c}^{\ ab}(x,x^{\prime})W(x,x^{\prime})+\tilde{\Sigma}_{c}^{\ ab},$ (78)
where $\tilde{\Theta}_{ab}$ and $\tilde{\Sigma}_{c}^{\ ab}$ are finite ambiguities that arise from the regularization procedure. We can decompose these ambiguities as
@@
$\tilde{\Sigma}_{c}^{\ ab}$ $={\Sigma^{M^{2}}}_{c}^{\ ab}+\Sigma_{c}^{\ ab},$ (80)
where $\Theta^{M^{2}}_{ab}$ and ${\Sigma^{M^{2}}}_{c}^{\ ab}$ are scale ambiguities, introduced to account for Eq. (43), while $\Theta_{ab}$ and $\Sigma_{c}^{\ ab}$ correspond to renormalization ambiguities. We explain the appearance of the term $g_{ab}v_{1}/(4\pi^{2})$ below.
We can explicitly compute $\omega(\tau_{ab})$ and $\omega(\sigma_{c}^{\ ab})$ by expanding $W$ using Eq. (56). Additionally, using Eqs. (108)-(114), we can bring these expectation values to the form
$\omega(\tau_{ab})$ $=\frac{1}{8\pi^{2}}\Bigg\{-\Bigg(w_{ab}-\frac{1}{2}g_{ab}w^{c}_{\ c}\Bigg)$
@@
$+\xi\mathring{G}_{ab}w-\frac{1}{2}m^{2}g_{ab}w+\xi[K^{c}_{\ c(a}\mathring{\nabla}_{b)}w$
$-K^{c}_{\ (ab)}\mathring{\nabla}_{c}w+(K^{d}_{\ (ab)}K^{c}_{\ cd}-K^{d}_{\ c(a}K^{c}_{\ b)d})w$
$-g_{ab}(K^{d\ c}_{\ [c}K^{e}_{\ e]d}w-K^{c\ d}_{\ d}\mathring{\nabla}_{c}w)]\Bigg\}+\tilde{\Theta}_{ab},$ (81)
@@
$+(g^{ab}K^{d}_{\ dc}-\delta^{a}_{\ c}K^{d\ b}_{\ d}+K_{c}^{\ ba}-K^{b\ a}_{\ c})w\Big\}$
$+\tilde{\Sigma}_{c}^{\ ab}.$ (82)
@@
$+\mathring{\nabla}^{a}K^{c}_{\ c[b}\mathring{\nabla}_{a]}w-\mathring{\nabla}^{a}K^{c}_{\ (ab)}\mathring{\nabla}_{c}$w
$+[K^{d}_{(ab)}K^{c}_{\ cd}-K^{d}_{\ c(a}K^{c}_{\ b)d}-\mathring{\nabla}_{b}K^{c}_{\ ca}]\mathring{\nabla}^{a}w$
$+[\mathring{\nabla}^{a}K^{d}_{(ab)}K^{c}_{\ cd}+K^{d}_{(ab)}\mathring{\nabla}^{a}K^{c}_{\ cd}$
$-\mathring{\nabla}^{a}K^{d}_{\ c(a}K^{c}_{\ b)d}-K^{d}_{\ c(a}\mathring{\nabla}^{a}K^{c}_{\ b)d}$
@@
$(K^{c}_{\ bd}\mathring{\nabla}_{a}+K^{c}_{\ ab}\mathring{\nabla}_{d}+\mathring{\nabla}_{a}K^{c}_{\ bd}+\mathring{\nabla}_{d}K^{c}_{\ ab})\omega(\sigma_{c}^{\ ad})$
$-(K^{d}_{\ ac}\mathring{\nabla}_{d}+\mathring{\nabla}_{d}K^{d}_{\ ac})\omega(\sigma_{b}^{\ ac})-\mathring{\nabla}_{b}K^{c}_{\ ad}\omega(\sigma_{c}^{\ ad})$
@@
$+(K^{c}_{\ bd}\mathring{\nabla}_{a}+K^{c}_{\ ab}\mathring{\nabla}_{d}+\mathring{\nabla}_{a}K^{c}_{\ bd}+\mathring{\nabla}_{d}K^{c}_{\ ab})\tilde{\Sigma}_{c}^{\ ad}$
$-(K^{d}_{\ ac}\mathring{\nabla}_{d}+\mathring{\nabla}_{d}K^{d}_{\ ac})\tilde{\Sigma}_{b}^{\ ac}-\mathring{\nabla}_{b}K^{c}_{\ ad}\tilde{\Sigma}_{c}^{\ ad}.$ (84)
@@
$(K^{c}_{\ bd}\mathring{\nabla}_{a}+K^{c}_{\ ab}\mathring{\nabla}_{d}+\mathring{\nabla}_{a}K^{c}_{\ bd}+\mathring{\nabla}_{d}K^{c}_{\ ab})\tilde{\Sigma}_{c}^{\ ad}$
$-(K^{d}_{\ ac}\mathring{\nabla}_{d}+\mathring{\nabla}_{d}K^{d}_{\ ac})\tilde{\Sigma}_{b}^{\ ac}-\mathring{\nabla}_{b}K^{c}_{\ ad}\tilde{\Sigma}_{c}^{\ ad}.$ (85)
Note that the term $g_{ab}v_{1}/(4\pi^{2})$ included in Eq. (79) cancels a similar term in Eq. (5.2). Hence, the left-hand side of Eq. (5.2) corresponds to the divergence of $\Theta^{M^{2}}_{ab}+\Theta_{ab}$; consequently, $\Theta^{M^{2}}_{ab}+\Theta_{ab}$ and ${\Sigma^{M^{2}}}_{c}^{\ ab}+\Sigma_{c}^{\ ab}$ satisfy an equation analogous to that in the third axiom. Moreover, $\omega(\tau_{ab})$ can be further simplified by substituting Eq. (60) into Eq. (81), yielding
@@
$+\xi\Big[K^{c}_{\ c(a}\mathring{\nabla}_{b)}w-K^{c}_{\ (ab)}\mathring{\nabla}_{c}w$
$+(K^{d}_{\ (ab)}K^{c}_{\ cd}-K^{d}_{\ c(a}K^{c}_{\ b)d})w+g_{ab}\mathring{\nabla}_{c}(K^{c\ d}_{\ d}w)\Big]$
@@
We proceed to construct $\tilde{\Theta}_{ab}$ and $\tilde{\Sigma}_{c}^{\ ab}$.

### v2

@@ 3.3.2 Conformal transformations
$g^{ab}\rightarrow\Omega^{-2}g^{ab},\quad K^{c}_{\phantom{c}ab}\rightarrow K^{c}_{\phantom{c}ab},\quad\Phi\rightarrow\Omega^{-1}\Phi,$
@@
$\delta S_{\text{M}}=\int\epsilon\,\tau^{a}_{\phantom{a}a}\,\sqrt{-g}\,\mathrm{d}^{4}x.$ (36)
@@
$\tau^{a}_{\phantom{a}a}=0.$ (37)
For the Klein–Gordon field under consideration in the special case where $m=0$ and $\xi=1/6$, and the fields are on-shell, this reduces to
$\tau^{a}_{\phantom{a}a}=\frac{1}{3}\mathring{\nabla}_{a}\big(K^{a\phantom{b}b}_{\phantom{a}b}\,\Phi^{2}\big),$ (38)
@@ 4 Hadamard bi-parametrix
To obtain well-defined operators associated with $\tau_{ab}$ and $\sigma_{c}^{\phantom{c}ab}$, which are quadratic in $\hat{\Phi}$, we employ point-splitting renormalization [38]. In this scheme, the central object is the two-point function
@@
We now insert Eq. (4) into Eq. (4) for $x\neq x^{\prime}$. From Eqs. (107) and (117) in Appendix A, we obtain
@@
where $\Delta$ is the Van Vleck–Morette determinant (see Appendix A). All arguments are omitted for brevity. By collecting the terms with the same $\sigma$-dependence, we can extract the following relations:
@@
From Eqs. (55) and (56), we can verify that $q_{a_{1}\cdots a_{p}}=q_{(a_{1}\cdots a_{p})}$, where $q$ stands for either $v_{n}$ or $w_{n}$. This symmetry, together with the exchange symmetry $x\leftrightarrow x^{\prime}$ of $V_{0}$ and $W$, allows us to express the coefficients with an odd number of indices in terms of those with an even number of indices. This is achieved by using Eqs. (108)-(115) given in Appendix A, explicitly implementing the exchange symmetry in Eqs. (55) and (56), and then taking the coincidence limit. The resulting identities are
@@
$w^{a}_{\phantom{a}a}$ $=(m^{2}+\xi\mathring{R})w-6v_{1}+2\xi(\mathring{\nabla}_{a}K^{a\phantom{b}b}_{\phantom{a}b}+K^{c\phantom{[a}a}_{\phantom{c}[a}K^{b}_{\phantom{b}b]c})w,$ (60)
@@
${v_{0}}^{a}_{\phantom{a}a}$ $=(m^{2}+\xi\mathring{R})v_{0}-4v_{1}+2\xi(\mathring{\nabla}_{a}K^{a\phantom{b}b}_{\phantom{a}b}+K^{c\phantom{[a}a}_{\phantom{c}[a}K^{b}_{\phantom{b}b]c})v_{0},$ (62)
@@
$-\frac{1}{180}\mathring{R}^{cd}\mathring{R}_{acbd}-\frac{1}{180}\mathring{R}_{a}^{\phantom{a}cde}\mathring{R}_{bcde},$ (67)
$v_{1}$ $=\frac{1}{8}m^{4}+\frac{1}{4}\xi m^{2}R-\frac{1}{24}m^{2}\mathring{R}-\frac{1}{24}\xi\square R+\frac{1}{120}\square\mathring{R}+\frac{1}{8}\Big(\xi R-\frac{1}{6}\mathring{R}\Big)^{2}-\frac{1}{720}\mathring{R}_{ab}\mathring{R}^{ab}+\frac{1}{720}\mathring{R}_{abcd}\mathring{R}^{abcd}.$ (68)
@@ 5 Semiclassical Einstein–Cartan theory
$\mathring{G}_{ab}+K^{d}_{\phantom{d}(ab)}K^{c}_{\phantom{c}cd}-K^{d}_{\phantom{d}c(a}K^{c}_{\phantom{c}b)d}$ $-K^{d\ \ c}_{\ [c}K^{e}_{\phantom{e}e]d}g_{ab}$
@@
$g^{ab}K^{d}_{\phantom{d}dc}-\delta^{a}_{\phantom{a}c}K^{d\ b}_{\ d}+K_{c}^{\phantom{c}ba}-K^{b\ a}_{\ c}$ $=16\pi\,\omega(\sigma_{c}^{\phantom{c}ab}),$ (70)
@@ 5.1 Axiomatic framework
    is smooth for any pair of test functions of compact support, $f_{1}$, $f_{2}$, then the differences $\omega_{1}(\tau_{ab})-\omega_{2}(\tau_{ab})$ and $\omega_{1}(\sigma_{c}^{\phantom{c}ab})-\omega_{2}(\sigma_{c}^{\phantom{c}ab})$ are also smooth.
@@
    The expectation values $\omega(\tau_{ab}(x))$ and $\omega(\sigma_{c}^{\phantom{c}ab}(x))$ are local with respect to the state of the Klein–Gordon field in the following sense: Let $(\mathcal{M},g_{ab})$ and $(\mathcal{M}^{\prime},g_{ab}^{\prime})$ be globally hyperbolic spacetimes with Cauchy surfaces $\Sigma$ and $\Sigma^{\prime}$, respectively. Let $\mathcal{O}\subset\mathcal{M}$ and $\mathcal{O}^{\prime}\subset\mathcal{M}^{\prime}$ be globally hyperbolic open neighborhoods of $x$ and $x^{\prime}$, with Cauchy surfaces $\mathcal{O}\cap\Sigma$ and $\mathcal{O}^{\prime}\cap\Sigma^{\prime}$, such that there exists an isometry between $\mathcal{O}$ and $\mathcal{O}^{\prime}$. Under this isometry, we can identify the Klein–Gordon subalgebras ${\mathcal{A}_{\text{KG}}}_{\mathcal{O}}\subset\mathcal{A}_{\text{KG}}$ for $\mathcal{M}$ and ${\mathcal{A}^{\prime}_{\text{KG}}}_{\mathcal{O}^{\prime}}\subset\mathcal{A}^{\prime}_{\text{KG}}$ for $\mathcal{M}^{\prime}$. If the restrictions of the Hadamard states coincide, i.e. $\omega|_{{\mathcal{A}_{\text{KG}}}_{\mathcal{O}}}=\omega^{\prime}|_{{\mathcal{A}^{\prime}_{\text{KG}}}_{\mathcal{O}^{\prime}}}$, then we require that $\omega(\tau_{ab}(x))=\omega^{\prime}(\tau_{ab}(x^{\prime}))$ and $\omega(\sigma_{c}^{\phantom{c}ab}(x))=\omega^{\prime}(\sigma_{c}^{\phantom{c}ab}(x^{\prime}))$.
@@
    $(K^{c}_{\phantom{c}bd}\mathring{\nabla}_{a}+K^{c}_{\phantom{c}ab}\mathring{\nabla}_{d}+\mathring{\nabla}_{a}K^{c}_{\phantom{c}bd}+\mathring{\nabla}_{d}K^{c}_{\phantom{c}ab})\,\omega(\sigma_{c}^{\phantom{c}ad})$
    $-(K^{d}_{\phantom{d}ac}\mathring{\nabla}_{d}+\mathring{\nabla}_{d}K^{d}_{\phantom{d}ac})\,\omega(\sigma_{b}^{\phantom{b}ac})-\mathring{\nabla}_{b}K^{c}_{\phantom{c}ad}\,\omega(\sigma_{c}^{\phantom{c}ad}).$ (72)
@@
    For a vacuum $\Omega$ in flat, torsionless spacetime, $\omega_{\Omega}(\tau_{ab})=0$ and $\omega_{\Omega}(\sigma_{c}^{\phantom{c}ab})=0$.
The first two axioms determine $\omega(\tau_{ab})$ and $\omega(\sigma_{c}^{\phantom{c}ab})$ up to local curvature terms, which cannot diverge in the flat-spacetime limit, as required by the fourth axiom. The third axiom is directly inspired by the (classical) relation given in Eq. (32). Moreover, the fourth axiom ensures the consistency of the semiclassical approach with quantum field theory in torsion-free flat spacetime. Equipped with these axioms, we now perform the regularization of the expectation values of the energy–momentum and spin–density operators.
@@ 5.2 Renormalization
$-K^{c}_{\phantom{c}(ab)}\mathring{\nabla}_{c}-K^{c}_{\phantom{c}(ab)}g_{c}^{\phantom{c}c^{\prime}}\mathring{\nabla}_{c^{\prime}}$
$+K^{d}_{\phantom{d}(ab)}K^{c}_{\phantom{c}cd}-K^{d}_{\phantom{d}c(a}K^{c}_{\phantom{c}b)d}$
$-g_{ab}(K^{d\phantom{[c}c}_{\phantom{d}[c}K^{e}_{\phantom{e}e]d}-K^{c\phantom{d}d}_{\phantom{c}d}\mathring{\nabla}_{c}-K^{c\phantom{d}d}_{\phantom{c}d}g_{c}^{\phantom{c}c^{\prime}}\mathring{\nabla}_{c^{\prime}})],$ (73)
@@
$+g^{ab}K^{d}_{\phantom{d}dc}-\delta^{a}_{\phantom{a}c}K^{d\phantom{d}b}_{\phantom{d}d}+K_{c}^{\phantom{c}ba}-K^{b\phantom{c}a}_{\phantom{b}c}),$ (74)
where unprimed indices refer to $x$, and primed indices to $x^{\prime}$. In addition, $g_{a}^{\ a^{\prime}}$ is the bitensor that implements parallel transport from $x$ to $x^{\prime}$, as described in Appendix A. Concretely, the expectation values we need to compute are
@@
$\omega(\sigma_{c}^{\phantom{c}ab})$ $=\lim_{x^{\prime}\rightarrow x}\hat{\sigma}_{c}^{\phantom{c}ab}(x,x^{\prime})[-iG^{\text{F}}(x,x^{\prime})].$ (76)
@@
$\omega(\sigma_{c}^{\phantom{c}ab})$ $=\lim_{x^{\prime}\rightarrow x}\hat{\sigma}_{c}^{\phantom{c}ab}(x,x^{\prime})[-i(G^{\text{F}}(x,x^{\prime})-H_{\ell}(x,x^{\prime}))]$
$=\frac{1}{8\pi^{2}}\lim_{x^{\prime}\rightarrow x}\hat{\sigma}_{c}^{\phantom{c}ab}(x,x^{\prime})W(x,x^{\prime})+\tilde{\Sigma}_{c}^{\phantom{c}ab},$ (78)
where $\tilde{\Theta}_{ab}$ and $\tilde{\Sigma}_{c}^{\phantom{c}ab}$ are finite ambiguities that arise from the regularization procedure. We can decompose these ambiguities as
@@
$\tilde{\Sigma}_{c}^{\phantom{c}ab}$ $={\Sigma^{M^{2}}}_{c}^{\phantom{c}ab}+\Sigma_{c}^{\phantom{c}ab},$ (80)
where $\Theta^{M^{2}}_{ab}$ and ${\Sigma^{M^{2}}}_{c}^{\phantom{c}ab}$ are scale ambiguities, introduced to account for Eq. (43), while $\Theta_{ab}$ and $\sigma_{c}^{\phantom{c}ab}$ correspond to renormalization ambiguities. We explain the appearance of the term $g_{ab}v_{1}/(4\pi^{2})$ below.
We can explicitly compute $\omega(\tau_{ab})$ and $\omega(\sigma_{c}^{\phantom{c}ab})$ by expanding $W$ using Eq. (56). Additionally, using Eqs. (108)-(114), we can bring these expectation values to the form
$\omega(\tau_{ab})$ $=\frac{1}{8\pi^{2}}\Bigg\{-\Bigg(w_{ab}-\frac{1}{2}g_{ab}w^{c}_{\phantom{c}c}\Bigg)$
@@
$+\xi\mathring{G}_{ab}w-\frac{1}{2}m^{2}g_{ab}w$
$+\xi[K^{c}_{\phantom{c}c(a}\mathring{\nabla}_{b)}w-K^{c}_{\phantom{c}(ab)}\mathring{\nabla}_{c}w$
$+(K^{d}_{\phantom{d}(ab)}K^{c}_{\phantom{c}cd}-K^{d}_{\phantom{d}c(a}K^{c}_{\phantom{c}b)d})w$
$-g_{ab}(K^{d\phantom{[c}c}_{\phantom{d}[c}K^{e}_{\phantom{e}e]d}w-K^{c\phantom{d}d}_{\phantom{c}d}\mathring{\nabla}_{c}w)]\Bigg\}+\tilde{\Theta}_{ab},$ (81)
@@
$+(g^{ab}K^{d}_{\phantom{d}dc}-\delta^{a}_{\phantom{a}c}K^{d\phantom{d}b}_{\phantom{d}d}+K_{c}^{\phantom{c}ba}-K^{b\phantom{c}a}_{\phantom{b}c})w\Big\}$
$+\tilde{\Sigma}_{c}^{\phantom{c}ab}.$ (82)
@@
$+\mathring{\nabla}^{a}K^{c}_{\phantom{c}c[b}\mathring{\nabla}_{a]}w-\mathring{\nabla}^{a}K^{c}_{\phantom{c}(ab)}\mathring{\nabla}_{c}$w
$+[K^{d}_{\phantom{d}(ab)}K^{c}_{\phantom{c}cd}-K^{d}_{\phantom{d}c(a}K^{c}_{\phantom{c}b)d}-\mathring{\nabla}_{b}K^{c}_{\phantom{c}ca}]\mathring{\nabla}^{a}w$
$+[\mathring{\nabla}^{a}K^{d}_{\phantom{d}(ab)}K^{c}_{\phantom{c}cd}+K^{d}_{\phantom{d}(ab)}\mathring{\nabla}^{a}K^{c}_{\phantom{c}cd}$
$-\mathring{\nabla}^{a}K^{d}_{\phantom{d}c(a}K^{c}_{\phantom{c}b)d}-K^{d}_{\phantom{d}c(a}\mathring{\nabla}^{a}K^{c}_{\phantom{c}b)d}$
@@
$(K^{c}_{\phantom{c}bd}\mathring{\nabla}_{a}+K^{c}_{\phantom{c}ab}\mathring{\nabla}_{d}+\mathring{\nabla}_{a}K^{c}_{\phantom{c}bd}+\mathring{\nabla}_{d}K^{c}_{\phantom{c}ab})\omega(\sigma_{c}^{\phantom{c}ad})$
$-(K^{d}_{\phantom{d}ac}\mathring{\nabla}_{d}+\mathring{\nabla}_{d}K^{d}_{\phantom{d}ac})\omega(\sigma_{b}^{\phantom{b}ac})-\mathring{\nabla}_{b}K^{c}_{\phantom{c}ad}\omega(\sigma_{c}^{\phantom{c}ad})$
@@
$+(K^{c}_{\phantom{c}bd}\mathring{\nabla}_{a}+K^{c}_{\phantom{c}ab}\mathring{\nabla}_{d}+\mathring{\nabla}_{a}K^{c}_{\phantom{c}bd}+\mathring{\nabla}_{d}K^{c}_{\phantom{c}ab})\tilde{\Sigma}_{c}^{\ ad}$
$-(K^{d}_{\phantom{d}ac}\mathring{\nabla}_{d}+\mathring{\nabla}_{d}K^{d}_{\phantom{d}ac})\tilde{\Sigma}_{b}^{\ ac}-\mathring{\nabla}_{b}K^{c}_{\phantom{c}ad}\tilde{\Sigma}_{c}^{\ ad}.$ (84)
@@
$(K^{c}_{\phantom{c}bd}\mathring{\nabla}_{a}+K^{c}_{\phantom{c}ab}\mathring{\nabla}_{d}+\mathring{\nabla}_{a}K^{c}_{\phantom{c}bd}+\mathring{\nabla}_{d}K^{c}_{\phantom{c}ab})\tilde{\Sigma}_{c}^{\phantom{c}ad}$
$-(K^{d}_{\phantom{d}ac}\mathring{\nabla}_{d}+\mathring{\nabla}_{d}K^{d}_{\phantom{d}ac})\tilde{\Sigma}_{b}^{\phantom{b}ac}-\mathring{\nabla}_{b}K^{c}_{\phantom{c}ad}\tilde{\Sigma}_{c}^{\phantom{c}ad}.$ (85)
Note that the term $g_{ab}v_{1}/(4\pi^{2})$ included in Eq. (79) cancels a similar term in Eq. (5.2). Hence, the left-hand side of Eq. (5.2) corresponds to the divergence of $\Theta^{M^{2}}_{ab}+\Theta_{ab}$; consequently, $\Theta^{M^{2}}_{ab}+\Theta_{ab}$ and ${\Sigma^{M^{2}}}_{c}^{\phantom{c}ab}+\Sigma_{c}^{\phantom{c}ab}$ satisfy an equation analogous to that in the third axiom. Moreover, $\omega(\tau_{ab})$ can be further simplified by substituting Eq. (60) into Eq. (81), yielding
@@
$+\xi\Big[K^{c}_{\phantom{c}c(a}\mathring{\nabla}_{b)}w-K^{c}_{\phantom{c}(ab)}\mathring{\nabla}_{c}w$
$+(K^{d}_{\phantom{d}(ab)}K^{c}_{\phantom{c}cd}-K^{d}_{\phantom{d}c(a}K^{c}_{\phantom{c}b)d})w+g_{ab}\mathring{\nabla}_{c}(K^{c\phantom{d}d}_{\phantom{c}d}w)\Big]$
@@
We proceed to construct $\tilde{\Theta}_{ab}$ and $\tilde{\Sigma}_{c}^{\phantom{c}ab}$.